OpenStudy (anonymous):

y=-4x+2 y=-3x+1 im looking for the plotted points ( , )

OpenStudy (anonymous):

@satellite73

OpenStudy (mattyboyy):

I think you need to separate the x and solve for y? I think and vise versa

OpenStudy (nincompoop):

when you have two equations like that, you are going to need to solve for the intersection of both equations

OpenStudy (anonymous):

Solve for y and x intercepts, to graph it fast ;P

OpenStudy (nincompoop):



OpenStudy (anonymous):

:o but how would I know where to graph them?

OpenStudy (anonymous):

my teacher said the anwser to this one is (1,-2) but I don't really understand how he figured that out.

OpenStudy (nincompoop):

don't you know how to graph linear equations? you don't even need to graph it really, you can solve it analytically use substitution or elimination method

OpenStudy (anonymous):

oooh kk

OpenStudy (anonymous):

i think i got it

OpenStudy (nincompoop):

do you know how to do that?

OpenStudy (anonymous):

yeah

OpenStudy (nincompoop):

okay, work out your solution on a paper then show it here so I can check if you did it right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-4x+2-3x+1 +3 +3 -1+2+1 x=2

OpenStudy (nincompoop):

what are you doing?

OpenStudy (anonymous):

isn't that how you do it?... im dumb. Should I just ask my teacher for help in the morning

OpenStudy (nincompoop):

that is not how you do it if you're trying to use the elimination method you will have to solve either the x or y first and in order for you to do that you will have to do the following: (1) \(y = -4x + 2 \) (2) \(y= -3x+1 \) multiply -1 to the first equation so you can eliminate the y first (1) \((y = -4x + 2 ) \times -1\) (2) \(y= -3x+1 \) this becomes: (1) \(-y = +4x - 2 \) (2) \(y= -3x+1 \) solve this one for me

OpenStudy (anonymous):

There's many ways to graph it, you can just use y =mx+b where m is the slope and b is the y intercept (where the line crosses the y - axis)...or you can solve for the intercepts then connect the dots... To find intercepts, for each equation, to find x intercept set y = 0 and solve for x, to find y - intercept set each equation x = 0 and solve for y.

OpenStudy (anonymous):

so im solving y=-3x+1 so its (y=-3x+1)x-1

OpenStudy (nincompoop):

no

OpenStudy (nincompoop):

(1) \(-y = +4x - 2 \) (2) \( y = -3x +1 \) SOLVE for this one first

OpenStudy (anonymous):

I'm not understanding. Thanks for trying to help ninc and batman. I'm just gonna go back to graphing it then just putting down where they interesect. It's easier for me that way.

OpenStudy (anonymous):

Nins weird, I hate his way

OpenStudy (nincompoop):

you are doing more work by doing that way, and you won't be able to figure out the intersection right away

OpenStudy (anonymous):

elimination is so gross

OpenStudy (anonymous):

At least i'll figure it out and be able to turn my homework in tomorrow morning XD

OpenStudy (nincompoop):

all I am asking you to do is SOLVE the 2 equations I set up for you it only involves subtraction

OpenStudy (nincompoop):

do you know how to subtract?

OpenStudy (anonymous):

1x + 1

OpenStudy (nincompoop):

can you show AN equation where is the equal sign?

OpenStudy (anonymous):

Haha, well basically what you have to do is add the equations that's why nin multiplied by - 1

OpenStudy (anonymous):

so the y cancels out and you can solve for x

OpenStudy (anonymous):

OH!

OpenStudy (anonymous):

4x - 3x = 1x 2- 1 = 1

OpenStudy (nincompoop):